OpenStudy (anonymous):

Help me! Write the equation of the line that is parallel to the line y = -3x + 12 and passes through the point (-1, 6).

OpenStudy (anonymous):

y = -3x + 3

OpenStudy (anonymous):

Write the equation of the line that is perpendicular to the line y = 2x + 2 and passes through the point (6, 3).

OpenStudy (anonymous):

y = -1/2x + 6

OpenStudy (anonymous):

Write the equation of the line that is parallel to the line 3x - y = -3 and passes through the point (4, -2).

OpenStudy (anonymous):

y = -3x + 10

OpenStudy (anonymous):

D:> That answer isn't there.... :c y = -one third x - 6 y = -one third x - 14 y = 3x - 6 y = 3x - 14

OpenStudy (anonymous):

Opps I mad a mistake hold

OpenStudy (anonymous):

*made

OpenStudy (anonymous):

It is y = 3x - 14

OpenStudy (anonymous):

Well I would like to help more but i have to study for science :c

OpenStudy (anonymous):

Doesn't matter you've been helping me a lot. Thanks. c:

OpenStudy (anonymous):

No problem :) good luck

OpenStudy (anonymous):

Right back at ya'!